# Single-Image based unsupervised joint segmentation and denoising

### v1

@@ 2.1 Convex Chan-Vese Formulation
In [12], Chan et al propose to relax the binary Chan-Vese segmentation problem (1) and let the desired solution $u(x)$ take values in $[0,1]$. The resulting convex energy is
[…]
@@
Though the model is convex it still suffers from difficulties in segmenting images where the piecewise constant assumption is not a relevant prior for the different regions in the image, or if the image is corrupted by severe noise. These issues are the main problems to solve in the current paper.
@@ 2.2 Self-supervised single-image based denoising
For a given noisy image $f=g+n$, self-supervised denoising methods are based on some variant of the self supervised loss
$\mathcal{L}_{f}(\theta)=\int_{\Omega}(\Phi_{\theta}(f)(x)-f(x))^{2}.$ (3)
Clearly, such a strategy cannot work without restricting the class of functions $\Phi_{\theta}$, since $\Phi_{\theta}=Id$ would be the minimiser and does not yield a denoised image. One strategy to overcome this problem is the method introduced in [39] where a generative model $\Phi_{\theta}$ is trained minimising (3). In this framework, the convolutional structure and early stopping prevents $\Phi_{\theta}$ to learn the fine image features (noise) to obtain a denoised image. Another strategy is linear filtering with a restriction on the filter [3, 24]. For example, a filter which is zero in its central position, and therefore not taking into account the information of this pixel but only the surrounding areas can be used to denoise an image minimising (3). Another type of method tries to minimise a slightly different functional. Motivated by the previous example, the authors of [26] introduce $N$ random binary masks $\mathcal{H}_{k}$ that delete information in the image. Training is then done using the loss function
$\mathcal{L}_{f}(\theta)=\frac{1}{N}\sum_{k=1}^{N}\int(\Phi_{\theta}(\mathcal{H}_{k}\cdot f)(x)-f(x))^{2}\cdot(1-\mathcal{H}_{k}).$ (4)
This training strategy also prevents the network from learning the identity operator. Although not directly minimising (3), we use a variant of this method named Noise2Fast [26]. This variant uses regular masks $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3},\mathcal{H}_{4}$, which consist of horizontal and vertical stripes on the even and odd image indices.
@@ 3 Proposed Joint Denoising and Segmentation
We now introduce our joint model, inspired by the observations described in Section 1.5. To control binary segmentation, we propose to train two denoising neural networks, each focusing on performing well in one of the regions to be segmented (cf. Figure 1 and 2). We denote these “experts” by $\Phi_{\theta^{F}}$ for the foreground, and $\Phi_{\theta^{B}}$ for the background. These experts are neural networks with parameters $\theta^{F}$ and $\theta^{B}$, that are trained with a modified denoising strategy. Let us mention that the model is presented in the case of two regions, but the extension to multi-class is straightforward, following for instance the framework in [19, 1, 30]
In Section 3.1, we present the proposed joint energy function designed for the combined denoising and segmentation process. This energy generalizes the convex Chan-Vese functional (2) with a data-fidelity term defined from the self-supervised denoising method . The optimisation scheme is performed in an alternating way, as presented in Section 3.2. We finally provide theoretical convergence results for our algorithm in Section 3.3.
@@ 3.1 Joint energy functional
In the following, we denote by $BV(\Omega)$ the space of all integrable functions $u:\Omega\rightarrow\mathbb{R}$ with bounded total variation $|u|_{\text{TV}}$, and consider the admissible set
$\mathbb{A}\coloneqq\{u\in BV(\Omega)\mid 0\leq u\leq 1\}.$
Further, let $i_{\mathbb{A}}:BV(\Omega)\rightarrow[0,\infty]$ denote the associated indicator function, which is 0 inside $\mathbb{A}$, and $\infty$ elsewhere. The parameters of the two denoising experts, $\Phi_{\theta^{F}}$ and $\Phi_{\theta^{B}}$ are denoted by $\boldsymbol{\theta}=(\theta^{F},\theta^{B})\in\mathbb{R}^{L\times L},$ and are respectively dedicated to the foreground and the background. These two experts are neural networks trained using the strategy proposed in [26]. We consider the joint model
[…]
Note that for fixed network parameters $\boldsymbol{\theta}$, the proposed energy is convex in $u$. Moreover, we can threshold the result and still have a global optima (see Theorem 3.2). Further we point out that in the case where the noise2fast training strategy is used, the energy functional for the denoising step is not exactly functional (5).
Figure 3 illustrates the idea behind the proposed segmentation model. For grayscale images, one can initialise the algorithm by thresholding image values. In more complex cases, a user can be asked to provide representative boxes for the background and foreground regions. Then, alternately, the denoising experts are trained on subsets of the two different segmented regions and the segmentations are updated. In practice, the data-fidelity term in (5) is updated given the denoising performance of the two experts $\Phi_{\theta^{F}}$ and $\Phi_{\theta^{B}}$. For fixed network parameters $\boldsymbol{\theta}$, the energy (5) is minimised. Repeating this procedure until a convergence criteria is met, we obtain the segmentation mask $u$, as well as the denoised image $g\approx u\odot\Phi_{\theta^{F}}+(1-u)\odot\Phi_{\theta^{B}}$.
Figure 3: The first image shows the given grayscale input image $f$, and user defined boxes representing rough foreground and background regions. The third image highlights pixels where the foreground expert denoiser performs better than the background one, while the last image is the segmentation result obtained by minimising the proposed energy (5).
@@
  • Constant Background: In case where the background is assumed constant, one could simply assume that $\Phi_{\theta^{B}}(f)=\theta^{B}\mathbbm{1}$, which corresponds to estimate a scalar value $\theta^{B}$ being the mean value of the given image inside the corresponding region as in the original Chan and Vese model.
  • Linear filter: In this case, the network is linear with respect to the network parameters $\theta^{B}$, more precisely, $\Phi_{\theta^{B}}(f)=\omega_{\theta^{B}}\ast f$, leading to a bi-convex energy functional (5). In our toy example in Figure 1, we have applied such a linear network consisting of one single filter of kernel size $15\times 15$.
  • Filtering of data fidelity term: When one of the regions is assumed to be constant and high noise levels are present, mean filtering improves the results. The data fidelity terms of energy (5) can then be replaced by $\int_{\Omega}\left[K_{\sigma}\ast\left(f-\phi_{\theta^{F}}(f)\right)\right]^{2}u$ and $\int_{\Omega}\left[K_{\sigma}\ast\left(f-\phi_{\theta^{B}}(f)\right)\right]^{2}(1-u)$, respectively, where $K_{\sigma}$ is a mean filter with kernel size $\sigma$. A similar approach has been done in [27], where a more robust version of the Chan-Vese model [11] has been proposed by introducing Gaussian convolution in the data fitting terms, in order to make the method robust to non homogeneous regions.
@@ 3.2 Joint optimisation
We propose to iteratively optimise problem (5) with an alternating procedure [15]. In case the denoising step does not exactly minimise energy (5), we actually alternate between minimising two slightly different functionals. For the sake of readability, this is not indicated in the notation. We start with the initialiseation of the segmentation mask $u.$ This is either achieved by thresholding for grayscale images, or as shown in Figure 3, manually choosing boxes representing the different regions to segment in the image. Then, based on the initial guess, the denoising expert(s) $\Phi_{\theta^{F}},$ and $\Phi_{\theta^{B}}$ are trained on the given initial masks. To this end, we use the ADAM optimiser [22] until convergence. As a next step, for fixed network parameters $\boldsymbol{\theta}$, we update the segmentation mask $u$. For fixed $\boldsymbol{\theta},$ the energy functional (5) is convex, and all the necessary assumptions for the application of the primal dual algorithm [10] are fulfilled. A more detailed description on the considered discrete schemes is provided in Section 4 (see Algorithm 2). These alternate steps are repeated as long as the decrease of energy (5) is greater than $p=15$ percent, which we empirically found to give a good compromise between computation speed and quality of the results.
The overall joint optimisation scheme is presented in Algorithm 1. A sketch of the alternating procedure is provided in Figure 4.
@@
  Initialise $u^{0}\leftarrow\boldsymbol{1}_{\{f>\epsilon\}}$ and $\boldsymbol{\theta}^{0}=\boldsymbol{\theta}_{0}$
  while $\mathcal{E}^{k}_{f,\lambda}(u^{k},\boldsymbol{\theta}^{k})/\mathcal{E}^{k-1}_{f,\lambda}(u^{k-1},\boldsymbol{\theta}^{k-1})\geq p\cdot\mathcal{E}^{k-1}_{f,\lambda}(u^{k-1},\boldsymbol{\theta}^{k-1})/\mathcal{E}^{k-2}_{f,\lambda}(u^{k-2},\boldsymbol{\theta}^{k-2})$ do
      $\boldsymbol{\boldsymbol{\theta}}^{k+1}\leftarrow\operatorname*{argmin}_{\boldsymbol{\theta}}\mathcal{E}^{k}_{f,\lambda}(u^{k+1},\boldsymbol{\theta})$with a few ADAM iterations for $\theta^{F}$ and Chan and Vese update for the background if $\Phi_{\theta^{B}}(f)=\theta^{B}\mathbbm{1}$}
     $u^{k+1}\leftarrow\operatorname*{argmin}_{u}\mathcal{E}^{k}_{f,\lambda}(u,\boldsymbol{\theta}^{k})$with Algorithm 2)}
  end while
@@
Figure 4: Alternating optimisation scheme. As a first step, regions are provided for the training of the two denoising experts using the strategy. These regions can be obtained by thresholding image values or by manually choosing boxes. The differences between the given noisy image $f$ and network outputs $\Phi_{\theta^{F}}(f)$ and $\Phi_{\theta^{B}}(f)$, are used in the subsequent segmentation step, minimising $\mathcal{E}_{\lambda,f}(\cdot,\boldsymbol{\theta})$ with Algorithm 2.
@@ 3.3 Theoretical Results
The proposed energy functional (5) is continuous and bounded from below. Therefore, for each $k\geq 0$, the following relations hold
$\mathcal{E}_{f,\lambda}(u^{(k)},\boldsymbol{\theta}^{(k+1)})$ $\leq\mathcal{E}_{f,\lambda}(u^{(k-1)},\boldsymbol{\theta}^{(k)})$
$\mathcal{E}_{f,\lambda}(u^{(k+1)},\boldsymbol{\theta}^{(k)})$ $\leq\mathcal{E}_{f,\lambda}(u^{(k)},\boldsymbol{\theta}^{(k-1)}).$
Hence, the generated sequence $\{\mathcal{E}_{f,\lambda}(u^{(k)},\boldsymbol{\theta}^{(k)})\}_{k\in\mathbb{N}}$ converges monotonically.
Theorem 3.1 (Convergence of Algorithm 1).
Assume that the level set $S^{0}=\{(u,\boldsymbol{\theta}):\mathcal{E}_{f,\lambda}(u,\boldsymbol{\theta})\leq\mathcal{E}_{f,\lambda}(u^{0},\boldsymbol{\theta}^{0})\}$ of $\mathcal{E}_{f,\lambda}$ defined in (5) is compact and that $\mathcal{E}_{f,\lambda}$ is continuous on $S^{0}$. Then, the sequence $\{(u^{k},\boldsymbol{\theta}^{k}\}$ generated by Algorithm 1 is defined and bounded. Moreover, every cluster point of $\{(u^{k},\boldsymbol{\theta}^{k})\}$ is a stationary point of $\mathcal{E}_{f,\lambda}$.
@@
This is a direct application of Theorem 4.1 in [38], using that (i) we only alternate between two variables $u$ and $\boldsymbol{\theta}$, (ii) the coupling between $u$ and $\boldsymbol{\theta}$ in $\mathcal{E}_{f,\lambda}$ is smooth. ∎
@@
The energy (5), which is convex for fixed network parameters $\boldsymbol{\theta}=(\theta^{F},\theta^{B})$ is a relaxation of the fully non-convex problem
[…]
where $\Sigma\subset\mathbb{R}^{2}$, and $\Omega\setminus\Sigma$ are the two regions of the given image $f(x),$ and Per$(\Sigma,\Omega)$ is the perimeter of the interface seperating these two regions.
Theorem 3.2 (Thresholding).
For any fixed $\boldsymbol{\theta}$, a global minimiser for the non-convex problem $\min_{\Sigma,\boldsymbol{\theta}}\mathcal{E}(\cdot,\boldsymbol{\theta})$ in (6) can be found by carrying out the minimisation of $\min_{u}\mathcal{E}_{f,\lambda}(\cdot,\boldsymbol{\theta})$, and then setting $\Sigma=\{x:u(x)\geq\tau\}$ for a.e. $\tau\in[0,1]$.
@@
The proof is similar to the one in [12](Theorem 2). The only difference is in the data fidelity term, where instead of the fixed constants $c_{1}$, and $c_{2}$, we look at fixed network outputs $\Phi_{\theta^{F}}(f)$, and $\Phi_{\theta^{B}}(f)$. As the problem is one-homogeneous in $u$, thanks to the co-area formula, we show that $\mathcal{E}_{f,\lambda}(u,\boldsymbol{\theta})=\int_{\epsilon}\mathcal{E}_{f,\lambda}(\mathbbm{1}_{u>\tau},\boldsymbol{\theta})$, so we can thus conclude that if $u$ is a minimiser of the energy (5) for fixed $\boldsymbol{\theta}$, then for a.e. $\tau\in[0,1]$ the set $\Sigma(\tau)$ has to be a minimiser of (6). ∎

### v2

@@ 2.1 Convex Chan-Vese Formulation
In [10], Chan et al propose to relax the binary Chan-Vese segmentation problem and let the desired solution $u$ take values in $[0,1]$. The resulting convex energy is
$\min_{0\leq u\leq 1}\int_{\Omega}\lvert\nabla u(x)\rvert+\lambda\left(\int_{\Omega}(c_{1}-f(x))^{2}u(x)\operatorname{d}\!x+\int_{\Omega}(c_{2}-f(x))^{2}(1-u(x))\operatorname{d}\!x\right).$ (2)
@@
Though the model is convex it still suffers from difficulties in segmenting images where the piecewise constant assumption is not a suitable prior for the different regions in the image, or if the image is corrupted by severe noise. These issues are the main problems to solve in the current paper by substituting $c_{1}$ and $c_{2}$ by more flexible neural network denoisers.
@@ 2.2 Self-supervised single-image based denoising
Before presenting the self-supervised loss, we introduce a very important property of denoising functions we are interested in.
Formulated in a very general way, for a given noisy image $f=g+n$, self-supervised denoising methods are based on some variant of the self supervised loss
$\mathcal{L}_{f}(\mathbf{D})=\int_{\Omega}(\mathbf{D}(f)(x)-f(x))^{2}\operatorname{d}\!x,\,\mathbf{D}\in\mathcal{D}$ (3)
minimised with an optimisation strategy, that excludes the identity function, $\mathbf{D}=\rm{Id}$, which would be the exact minimiser but does not yield a denoised image.
One strategy to overcome this problem is the method introduced in [32] where a generative model $\mathbf{D}$ for minimising (3) is combined with early stopping. In this framework, the convolutional structure and early stopping prevents $\mathbf{D}$ to learn the fine image features (noise) to obtain a denoised image. Another strategy is linear filtering with a restriction on the filter [3, 17]. For example, a filter which is zero in its central position, and therefore not taking into account the information of this pixel but only the surrounding areas can be used to denoise an image minimising (3). Another type of method tries to minimise a slightly different functional. Motivated by the previous example, the authors of [19] introduce $N$ random binary masks $\mathcal{H}_{k}$ that delete information in the image. Training is then done using the loss function
$\mathcal{L}_{f}(\mathbf{D})=\frac{1}{N}\sum_{k=1}^{N}\int_{\Omega}(\-\mathbf{D}(\mathcal{H}_{k}\cdot f)(x)-f(x))^{2}\cdot(1-\mathcal{H}_{k})\operatorname{d}\!x.$ (4)
This training strategy prevents the network from learning the identity operator. In this work, we use a variant of Noise2Fast [19], which is numerically cheaper than Noise2Void or Noise2Self, and where $\mathcal{H}_{k},k=1,\dots,N$, consist of even and odd pixels of the image, respectively.
@@ 3 Proposed Joint Denoising and Segmentation
We now introduce our joint model, inspired by the observations described in Section 1.1. To control binary segmentation, we propose to train two denoising neural networks, each focusing on performing well in one of the regions to be segmented (cf. Figure 1 and 2). We denote these “experts” by $\mathbf{D}_{F}$ for the foreground, and $\mathbf{D}_{B}$ for the background. These experts are neural networks with parameters $F$ and $B$, that are trained with a modified denoising strategy. Let us mention that the model is presented in the case of two regions, but the extension to multi-class is straightforward, following for instance the framework in [14, 1, 25]
@@ 3.1 Joint energy functional
In the following, we denote by $\mathrm{BV}(\Omega)$ the space of all integrable functions $u:\Omega\rightarrow\mathbb{R}$ with bounded total variation $|u|_{\text{TV}}$, and consider the admissible set
$\mathbb{A}\coloneqq\{u\in\mathrm{BV}(\Omega)\mid 0\leq u\leq 1\}.$
Further, let $i_{\mathbb{A}}:\mathrm{BV}(\Omega)\rightarrow[0,\infty]$ denote the associated indicator function, which is 0 inside $\mathbb{A}$, and $\infty$ elsewhere. The parameters of the two denoising experts, $\mathbf{D}_{F}$ and $\mathbf{D}_{B}$ are denoted by $\boldsymbol{W}=(F,B)\in\mathbb{R}^{L_{1}\times L_{2}},$ and are respectively dedicated to the foreground and the background. Note that the number of parameters for $\mathbf{D}_{F}$ and $\mathbf{D}_{B}$ do not necessarily have to be the same. These two experts are neural networks trained using the strategy proposed in [19]. We consider the joint model
$\begin{split}\mathcal{E}_{f,\lambda}(u,\mathbf{D}_{F},\mathbf{D}_{B})=i_{\mathbb{A}}(u)+\lambda\lvert u\rvert_{\text{TV}}&+\int_{\Omega}\left(f(x)-\mathbf{D}_{F}(f)(x)\right)^{2}u(x)\operatorname{d}\!x\\ &+\int_{\Omega}\left(f(x)-\mathbf{D}_{B}(f)(x)\right)^{2}(1-u(x))\operatorname{d}\!x\,.\end{split}$ (5)
Note that for fixed network parameters $\boldsymbol{W}$, the proposed energy is convex in $u$. Moreover, we can threshold the result and still have a global optimum (see Theorem 4).
Figure 3 illustrates the idea behind the proposed segmentation model. For greyscale images, one can initialise the algorithm by thresholding image values. In more complex cases, a user can be asked to provide representative boxes for the background and foreground regions. Then, alternately, the denoising experts are trained on subsets of the two different segmented regions and the segmentations are updated. In practice, the data fidelity term in (5) is updated given the denoising performance of the two experts $\mathbf{D}_{F}$ and $\mathbf{D}_{B}$. For fixed network parameters $\boldsymbol{W}$, the energy (5) is minimised. Repeating this procedure until a convergence criteria is met, we obtain the segmentation mask $u$, as well as the denoised image $g\approx u\odot\mathbf{D}_{F}(f)+(1-u)\odot\mathbf{D}_{B}(f)$.
Figure 3: The first image shows the given greyscale input image $f$, and user defined boxes representing rough foreground and background regions. The third image highlights pixels where the foreground expert denoiser performs better than the background one, while the last image is the segmentation result obtained by minimising the proposed energy (5).
@@
  • Constant Background: In case where the background is assumed constant, one could simply assume that $\mathbf{D}_{B}(f)=B\odot\mathbbm{1}$, which corresponds to estimate a scalar value $B$ being the mean value of the given image inside the corresponding region as in the original Chan and Vese model.
  • Linear filter: In this case, the network is linear with respect to the network parameters $B$, more precisely, $\mathbf{D}_{B}(f)=\omega_{B}\ast f$, leading to a bi-convex energy functional (5). In our toy example in Figure 1, we have applied such a linear network consisting of one single filter of kernel size $15\times 15$.
  • Filtering of data fidelity term: When one of the regions is assumed to be constant and high noise levels are present, mean filtering improves the results. The data fidelity terms of energy (5) can then be replaced by $\int_{\Omega}\left[K_{\sigma}\ast\left(f-\mathbf{D}_{F}(f)\right)\right]^{2}u$ and $\int_{\Omega}\left[K_{\sigma}\ast\left(f-\mathbf{D}_{B}(f)\right)\right]^{2}(1-u)$, respectively, where $K_{\sigma}$ is a mean filter with kernel size $\sigma$. A similar approach has been done in [20], where a more robust version of the Chan-Vese model [9] has been proposed.
@@ 3.2 Joint optimisation
We propose to iteratively optimise problem (5) with an alternating procedure [12]. In case the denoising step does not exactly minimise energy (5), we actually alternate between minimising two slightly different functionals. For the sake of readability, this is not indicated in the notation. We start with the initialisation of the segmentation mask $u.$ This is either achieved by thresholding for greyscale images, or as shown in Figure 3, manually choosing boxes representing the different regions to segment in the image. Then, based on the initial guess, the denoising expert(s) $\mathbf{D}_{F},$ and $\mathbf{D}_{B}$ are trained on the given initial masks. To this end, we use the ADAM optimiser [15] until convergence. As a next step, for fixed network parameters $\boldsymbol{W}=(F,B)$, we update the segmentation mask $u$. For fixed $\boldsymbol{W},$ the energy functional (5) is convex, and all the necessary assumptions for the application of the primal dual algorithm [8] are fulfilled. A more detailed description on the considered discrete schemes is provided in Section 4 (see Algorithm 2). These alternate steps are repeated as long as the decrease of energy (5) is greater than $15$ percent (i.e. $p=0.15)$, which we empirically found to give a good compromise between computation speed and quality of the results.
The overall joint optimisation scheme is presented in Algorithm 1. For both denoising experts, we use the same network architecture, namely a convolutional neural network consisting of four sequential instances of convolutional layers, each of them consisting of 64 channels. The last layer is again a convolution with a $1\times 1$ kernel size, and the final output is obtained by applying sigmoid activation function. In each denoising update step, we use the same strategy as proposed in [19], namely after each training epoch, the MSE between the noisy input image and “denoised” network output is computed. If this metric does not change much for 100 epochs, training is terminated, and the average of the last one hundred validation images is the final denoised output. A sketch of the alternating procedure is provided in Figure 4.
  Input: noisy input image $f$
  Initialisation: $u^{0}\leftarrow\boldsymbol{1}_{\{f>\epsilon\}}$, $\boldsymbol{W}^{0}=\boldsymbol{W}_{0}$, choose $\lambda>0$ and set $p\leftarrow 0.15$
  while $\mathcal{E}^{k}_{f,\lambda}(u^{k},\boldsymbol{W}^{k})/\mathcal{E}^{k-1}_{f,\lambda}(u^{k-1},\boldsymbol{W}^{k-1})\geq p\cdot\mathcal{E}^{k-1}_{f,\lambda}(u^{k-1},\boldsymbol{W}^{k-1})/\mathcal{E}^{k-2}_{f,\lambda}(u^{k-2},\boldsymbol{W}^{k-2})$ do
      $\boldsymbol{\boldsymbol{W}}^{k+1}\leftarrow\operatorname*{argmin}_{\boldsymbol{W}}\mathcal{E}^{k}_{f,\lambda}(u^{k+1},\boldsymbol{W})$with a few ADAM iterations for $F$ and Chan and Vese update for the background if $\mathbf{D}_{B}(f)=\mathbf{D}_{B}\mathbbm{1}$}
     $u^{k+1}\leftarrow\operatorname*{argmin}_{u}\mathcal{E}^{k}_{f,\lambda}(u,\boldsymbol{W}^{k})$with Algorithm 2)}
  end while
@@
Figure 4: Alternating optimisation scheme. As a first step, regions are provided for the training of the two denoising experts using the strategy. These regions can be obtained by thresholding image values or by manually choosing boxes. The differences between the given noisy image $f$ and network outputs $\mathbf{D}_{F}(f)$ and $\mathbf{D}_{B}(f)$, are used in the subsequent segmentation step, minimising $\mathcal{E}_{\lambda,f}(\cdot,\boldsymbol{W})$ with Algorithm 2.
@@ 3.3 Theoretical Results
The proposed energy functional (5) is continuous and bounded from below. Further, for each $k>0$, the following relations hold
$\mathcal{E}_{f,\lambda}(u^{(k)},\boldsymbol{W}^{(k+1)})$ $\leq\mathcal{E}_{f,\lambda}(u^{(k)},\boldsymbol{W}^{(k)})$
$\mathcal{E}_{f,\lambda}(u^{(k+1)},\boldsymbol{W}^{(k+1)})$ $\leq\mathcal{E}_{f,\lambda}(u^{(k)},\boldsymbol{W}^{(k)}).$
Hence, the generated sequence $\{\mathcal{E}_{f,\lambda}(u^{(k)},\boldsymbol{W}^{(k)})\}_{k\in\mathbb{N}}$ converges monotonically.
Theorem 3 (Convergence of Algorithm 1).
Assume that the level set $S^{0}=\{(u,\boldsymbol{W}):\mathcal{E}_{f,\lambda}(u,\boldsymbol{W})\leq\mathcal{E}_{f,\lambda}(u^{0},\boldsymbol{W}^{0})\}$ of $\mathcal{E}_{f,\lambda}$ defined in (5) is compact and that $\mathcal{E}_{f,\lambda}$ is continuous on $S^{0}$. Then, the sequence $\{(u^{k},\boldsymbol{W}^{k}\}$ generated by Algorithm 1 is defined and bounded. Moreover, every cluster point of $\{(u^{k},\boldsymbol{W}^{k})\}$ is a stationary point of $\mathcal{E}_{f,\lambda}$.
@@
This is a direct application of Theorem 4.1 in [31], using that (i) we only alternate between two variables $u$ and $\boldsymbol{W}$, (ii) the coupling between $u$ and $\boldsymbol{W}$ in $\mathcal{E}_{f,\lambda}$ is smooth. ∎
@@
The energy (5), which is convex for fixed network parameters $\boldsymbol{W}=(F,B)$ is a relaxation of the fully non-convex problem
$\mathcal{E}(\Sigma,\boldsymbol{W})=\text{Per}(\Sigma,\Omega)+\int_{\Sigma}(f-\mathbf{D}_{F}(f))^{2}\operatorname{d}\!x+\int_{\Omega\setminus\Sigma}(f-\mathbf{D}_{B}(f))^{2}\operatorname{d}\!x,$ (6)
where $\Sigma\subset\mathbb{R}^{2}$, and $\Omega\setminus\Sigma$ are the two regions of the given image $f(x),$ and Per$(\Sigma,\Omega)$ is the perimeter of the interface separating these two regions.
Theorem 4 (Thresholding).
For any fixed $\boldsymbol{W}$, a global minimiser for the non-convex problem $\min_{\Sigma,\boldsymbol{W}}\mathcal{E}(\cdot,\boldsymbol{W})$ in (6) can be found by carrying out the minimisation of $\min_{u}\mathcal{E}_{f,\lambda}(\cdot,\boldsymbol{W})$, and then setting $\Sigma(\tau)=\{x:u(x)\geq\tau\}$ for a.e. $\tau\in[0,1]$.
@@
The proof is similar to the one in [10](Theorem 2). The only difference is in the data fidelity term, where instead of the fixed constants $c_{1}$, and $c_{2}$, we look at fixed network outputs $\mathbf{D}_{F}(f)$, and $\mathbf{D}_{B}(f)$. By rewriting the $\lvert u\rvert_{\text{TV}}$-term with the co-area formula we show that $\mathcal{E}_{f,\lambda}(u,\boldsymbol{W})=\int_{0}^{1}\mathcal{E}(\Sigma(\tau),\boldsymbol{W})-C\operatorname{d}\!\tau$, with $C=\int_{\Omega}(f-\mathbf{D}_{F}(f))^{2}\operatorname{d}\!x$ being independent of $u$. Thus, we conclude that if $u$ is a minimiser of the energy (5) for fixed $\boldsymbol{W}$, then for a.e. $\tau\in[0,1]$ the set $\Sigma(\tau)$ has to be a minimiser of (6). ∎